OpenStudy (johnnyjames84543553):

What is the coefficient of the term x7y in the expansion of (x + y)8? a.1 b.8 c.28 d.56

OpenStudy (sshayer):

\[T _{r+1}=C_{r}{8}x ^{8-r}y ^{r}\] put r=1\[T _{9}=C _{1}^{8}x^7y^1=\frac{ 8 }{ 1 }x^7y=8x^7y\] coefficient of x^7y=?